Compute the missing frequencies in the frequency distribution table given below which is partly produced based on 100 observations for which the mean was found to 1.86.
Accidents 0 1 2 3 4 5
Frequency 36 ? ? 15 5 5

Answer by stanbon(48568)

Compute the missing frequencies in the frequency distribution table given below which is partly produced based on 100 observations for which the mean was found to 1.86.
Accidents .......0 1 2 3 4 5
Frequency 36 ? ? 15 5 5
--------------------
The total frequendy is 100
Let # under "1" be "x"
Then # under "2" is 100-x-61="39-x"
If the mean is 1.86, (sum of accident # * freq)/100 = 1.86
sum of acc # * freq = 0*36+1*x+2*(39-x)+3*15+4*5+5*5=x+78-2x+90
= -x+168
So, (168-x)= 186
That makes x a negative number.
Something is wrong in the data.
--------
Chers,
Stan H.